# Synthetic Prior for Few-Shot Drivable Head Avatar Inversion

### v2

@@ 1 Introduction
Another recent body of work addresses the problem of building 3D head avatars from one or few input images, [77, 8, 11]. However, their rendering quality and fidelity are typically lower than those of methods trained on large datasets (e.g., [50, 88]). To improve quality, some methods [81, 75, 76] first learn a multi-identity head model that is used as prior when optimizing for the personalized avatar. Training these head priors requires a large-scale multi-view image dataset that is expensive and time-consuming to capture. Moreover, managing real data under protection laws like GDPR is cumbersome for experimentation and maintenance, as users must periodically (e.g., every 30 days) delete all dataset derivatives and trained models, allowing dataset participants to be removed from both if needed. Alternatively, the FFHQ dataset [29] may be employed, with 4D GAN-based methods [80, 61, 13] constructing an inversion prior from it. However, these approaches tend to exhibit artifacts during novel view synthesis and struggle with preserving identity. In summary, the expressive power of this prior is strongly influenced by several factors: the diversity of the training data (e.g., ethnicity, age, facial features, expressions), the multi-view capture hardware setup (i.e., lighting, view-density, calibration quality, frame-rate), and the quality of the data pre-processing (e.g., mesh tracking, background masking).
@@
In contrast to the previous work that focuses on expensive and cumbersome real data, we overcome these limitations and propose SynShot, a new method that builds a prior solely on synthetic data and adapts to a real test subject requiring only a few input images. Building on the success of ML models trained on synthetic data for tasks like 3D face regression [57], 2D landmark prediction [70], rigid face alignment [3], and few-shot head reconstruction [68, 78, 6], SynShot is trained solely on a large synthetic dataset generated from 3DMM samples and diverse assets. Synthetic data offers complete control over dataset creation to meet size and diversity needs for training an expressive head prior, eliminating the need for costly capture hardware and addressing privacy concerns with real subjects. The benefits brought by synthetic data come at the cost of having to handle the domain gap between the trained head prior and real images captured “in the wild”. To effectively bridge this gap, we first fit the synthetic prior to real images and then fine-tune the prior weights to the real data using the pivotal tuning strategy proposed in [52]. With as few as three input images, SynShot reconstructs a photorealistic head avatar that generalizes to novel expressions and camera viewpoints (Fig. 1). The results show that our method outperforms state-of-the-art personalized monocular methods [85, 71, 59] trained on thousands of images each. Our method represents head avatars using 3D Gaussian primitives [30], where Gaussian parameters are generated by a convolutional architecture in the UV (texture) plane, similar to related work [41, 88, 55].
@@
    A novel generative method based on a convolutional encoder-decoder architecture that is trained on extensive synthetic data only to produce controllable 3D head avatars.
@@ 2 Related Work
Recent methods by Xu et al. [76] are conceptually similar to HeadNeRF and MofaNeRF; however, instead of embedding NeRF [45] on a mesh, they employ 3DGS [30]. GPHM [76] uses a series of MLPs to generate Gaussian primitives attached to a parametric model, enabling expression control and inversion, though it conditions only the avatar’s shape. In contrast, our approach incorporates both shape and color information. The follow-up work GPHMv2 [75] extends GPHM with a dynamic module for improved reenactment control and a larger dataset, further enhancing quality. HeadGAP [81] also models avatars using MLPs, utilizing part-based features and additional color conditioning to improve quality. While these methods embed primitives directly on the mesh surface, our approach explicitly learns the primitive parameters by modeling their distribution via a VQ-VAE [66], eliminating the need for a guiding mesh during the test time as the shape is captured within our latent space.
@@
Volumetric primitives, combined with multi-view training, are highly effective for modeling human heads [88, 55, 74, 23, 32, 28, 63, 62] as they capture intricate details like hair and subsurface scattering [56], outperforming traditional textured meshes. VolTeMorph [21] embeds a NeRF within tetrahedral cages that guide volumetric deformation. Qian et al. [50] attach Gaussian primitives to 3DMM triangles, whose local rotations and stretch deform the Gaussians without requiring neural networks. Xu et al. [74] and Giebenhain et al. [23] extend that work to further predict corrective fields over the Gaussians; rather than colors, they splat features that are translated into color by an image-space CNN [65]. Lombardi et al. [42] position 3D voxels with RGB and opacity values at the vertices of a head mesh, using ray tracing for volumetric integration. Saito et al. [55] improve quality by replacing voxel primitives with 3D Gaussians and applying rasterization. Our VQ-GAN training aligns with these principles for few-shot capture, as we supervise the process using a hybrid mesh-primitive approach to model the generative distribution.
@@
Monocular methods often rely on a strong 3DMM prior, as recovering a 3D shape from a 2D image is an inherently under-constrained problem. Face2Face [64] was a seminal work that enabled real-time reconstruction and animation of a parametric model. However, it lacks detailed hair representation and relies on low-frequency PCA texture models, which significantly affects quality. This limitation has led to the rise of neural avatars based on Neural Radiance Fields (NeRF) [67, 77, 73, 82, 72, 10, 82, 83, 18, 20, 19] and later on 3D Gaussian Splatting (3DGS) primitives [71, 59, 33, 9]. INSTA [85] applies triangle deformation gradients [60] to each NeRF sample based on proximity to the nearest triangle, enabling avatar animation. This approach has been extended to 3D Gaussian Splatting (3DGS)-based avatars by methods like Flash Avatar [71] and Splatting Avatar [59]. As Qian et al. [50], these methods extend a 3DMM with additional MLPs to model non-rigid deformations and fine details like wrinkles. Unlike SynShot, these monocular methods do not generalize well to novel views and expressions. Moreover, they require three orders of magnitude more real data to create a single avatar. Our method overcomes these limitations by leveraging a synthetic prior during avatar inversion, achieving high-quality results with only a few shots.
@@ 3 Method
This section describes SynShot, how we train the synthetic prior to generate drivable 3D Gaussian head avatars, and how we use it to do a few-shot reconstruction of a realistic head avatar from few images.
Figure 3: Latent space interpolation. Linearly interpolating $z_{\mathbf{q}_{\text{id}}}$ and $z_{\mathbf{q}_{\text{expr}}}$ between the leftmost and rightmost avatars demonstrates that our latent manifold exhibits smooth transitions in both expression and identity.
@@ 3.1 Preliminaries
Using a 3DMM, we represent a base 3D mesh as $\mathbf{S}=\bar{\mathbf{S}}+\boldsymbol{\delta}\mathbf{B}_{id}+\boldsymbol{\gamma}\mathbf{B}_{expr}$, where $\bar{\mathbf{S}}$ is the average shape, $\mathbf{B}_{(id,expr)}$ are the bases for identity and expression, and $\boldsymbol{\delta}$, $\boldsymbol{\gamma}$ denote the corresponding coefficients. Additionally, we use linear blend skinning (LBS) for head rotation around the neck with pose corrective offsets, and to rotate the eyeballs.
The head avatar is rendered via 3D Gaussian splatting, using the CUDA implementation of 3DGS [30]. The rasterizer is defined as $\mathcal{R}(\boldsymbol{G},\mathbf{K})\rightarrow\bar{\mathcal{I}}$, for a camera $\mathbf{K}$ and a set of $n$ 3D Gaussians $\boldsymbol{G}\in\mathbb{R}^{n\times(11+16\times 3)}:=\{\boldsymbol{\phi},\boldsymbol{\theta},\boldsymbol{\sigma},\boldsymbol{\alpha},\boldsymbol{h}\}$, with position $\boldsymbol{\phi}\in\mathbb{R}^{n\times 3}$, rotation $\boldsymbol{\theta}\in\mathbb{R}^{n\times 3\times 3}$, scale $\boldsymbol{\sigma}\in\mathbb{R}^{n\times 3}$, opacity $\boldsymbol{\alpha}\in\mathbb{R}^{n}$, and the (third-degree) spherical harmonics parameters $\boldsymbol{h}\in\mathbb{R}^{n\times 16\times 3}$, where $n$ is the number of Gaussians. Please refer to Kerbl et al. [30] for more details.
@@ 3.2 Gaussian Prior Model
A common limitation of using CNNs to regress Gaussian maps is the fixed output resolution, which ties the number of primitives to the output dimensions. This restriction can significantly limit the quality of the reconstructed avatar (see Table 1). To address this issue, we use a part-based densification mechanism. Similar to Kirschstein et al. [33], we use bilinear sampling, $\mathcal{B}(\cdot,u,v)$ to sample the output of the decoders at UV-positions $(u,v)$. As different head regions $r\in\{face,hair\}$ have varying requirements for the density of Gaussian primitives, we bilinearly sample separate parameter maps for the face and scalp region, rather than a single joint map.
Thus, per-part map sampling acts as adaptive primitive densification for the individual regions to improve visual quality (Table 1). While we use a higher number of primitives only for the hair and face regions, it can also be applied to the lower half of the face, for example, to model full beards.
@@ 3.3 Few-shot Avatar Reconstruction
To make the problem tractable, we employ a few heuristics to aid the optimization. These include early stopping with a warmup phase and an exponential moving average on the loss to determine the stopping criteria. Additionally, we scale the number of optimization steps based on the number of training frames, using a constant factor of 10 to increase the likelihood that each sample is seen at least once. As a training objective, in addition to our photometric term $\mathcal{L}_{\text{color}}$ (Eq. 5), we follow Lattas et al. [39] and, based on ArcFace [12], define two additional objectives: $\mathcal{L}_{\text{id}}$ and $\mathcal{L}_{\text{arc}}$. The final inversion loss is equal to $\mathcal{L}=\mathcal{L}_{color}+\mathcal{L}_{arc}+\mathcal{L}_{id}$. For a number of views, up to 20, the optimization takes less than 10 minutes on a single Nvidia H100 which is comparable to INSTA [85]. The training time increases with the number of frames as we scale the number of iterations accordingly.
@@ 3.4 Synthetic Dataset
Our dataset consists of approximately 2,000 unique identities, which we render with resolution $768\times 512$ using Blender (Cycles), following Wood et al. [70]. We positioned fourteen cameras in front of the subject and an additional fourteen cameras sampled from the upper hemisphere, centered on the scene. We randomly assign assets such as hairstyles and beards to these avatars. Additionally, we utilize high-quality face textures which are randomly distributed among the samples. By combining different shapes and appearances, we augment the set of identities, following practices in synthetic data [70] and 3D face reconstruction [39, 46, 14]. To incorporate tracked expressions from multi-view setups, we propagate them to the avatars during sequence rendering. We additionally compute a hair proxy from strands by voxelizing and fitting it to the scalp region; we apply the same approach for beards. Using a neutral mesh and its hair proxy, we backproject the images onto the texture map. During test time, we use a 3DMM regressor and the input images to extract a texture, which is then used as an initialization for our method. In total, our dataset comprises 14 million images for building the prior. Figure 21 shows random samples from the dataset.
@@ 4 Results
We compare SynShot to two different types of methods, state-of-the-art personalized monocular methods, and inversion-based general methods. The personalized monocular methods are controlled by FLAME [40] meshes and include INSTA [85], Flash Avatar [71], and SplattingAvatar [59]. For monocular methods, we used an ensemble of four datasets [25, 88, 82, 19] processed using the face tracker from Zielonka et al. [86]. SplattingAvatar follows the approach of Zheng et al. [82] and uses the monocular 3D face regressor DECA [17] for tracking. In our experiments, we adopted a similar approach, employing an in-house regressor, similar to DECA, to estimate 3DMM expression and pose parameters. While these methods produce photorealistic avatars, they struggle with generalization to novel views and poses (see Figure 16).
Figure 7: Novel view evaluation of long hair and beard inversion using only three input images demonstrates the strong generalization capability of SynShot, which accurately models both attributes. See supp. mat. for more results.
For inversion-based methods, we compare PanoHead [2], HeadNeRF [26], and MofaNeRF [84] (see supplemental material for details). We also compare SynShot to newer 4D GAN-based methods offering animation, including Portrait4D [13], Next3D [61], and InvertAvatar [80] (Figures 10 and 8). Moreover, Figure 7 presents a novel view evaluation of challenging long hair and beard inversion, demonstrating strong generalization capabilities of SynShot. We used three input images for all the inversion experiments which can be seen in the supplemental materials.
@@
To measure the performance of SynShot without introducing bias, we selected training frames from $\{F_{n}\}_{n=1}^{16}=\{1,\dots,987\}$, where $F_{n}$ denotes the Fibonacci sequence. For all experiments, we use progressive farthest point sampling [49] in the 3DMM expression space to select a specified number of frames from the training set. The self-reenactment sequences were evaluated using photometric metrics (LPIPS and SSIM) on the last 600 frames from the INSTA dataset [85].
Figure 8: The GAN-based cross-reenactment comparison shows that our method significantly better reconstructs the target subject’s appearance (identity) and remains faithful to the source subject’s head poses and expressions, whereas the other methods suffer from artifacts. Inversion input images can be found in supplemental materials.
@@
Our combined dataset consists of eleven monocular sequences ($512\times 512$ resolution), many of which are in-the-wild videos with very limited head motion, resulting in a low error as the test sequences closely resemble the training data, leaving limited room to assess diversity. To address this and accurately measure the effective error, we trained each method on a varying number of frames, corresponding to frames used in our inversion pipeline. The reconstruction error is evaluated on 600 test frames. Figure 9 demonstrates the effectiveness of our inversion, particularly with up to 233 training frames. Due to the lack of a strong prior, monocular methods fail in low training frame regimes, and, even with larger training datasets, they do not perform well and produce artifacts. Please note that to benefit from an increased number of input frames, i.e., to ground the avatar reconstruction more on the input than the synthetic prior, it requires an increased number of optimization iterations during pivotal fine-tuning. The number of iteration steps affects the reconstruction metrics, causing LPIPS to vary non-monotonically. Refer to the supp. mat. for qualitative evaluation.
@@
Figure 10: The GAN-based self-reenactment comparison again shows that SynShot better captures identity and synthesizes novel views, proving its usefulness as a synthetic prior and for pivotal fine-tuning in inversion. LPIPS scores: Ours (0.0236), InvertAvatar (0.0962), Portrait4D (0.0843), and Next3D (0.2274). Inversion input images can be found in supplemental materials.
@@
We compared SynShot to three animatable GAN-based methods. For our method and InvertAvatar [80], we used three input images, whereas Portrait4D [13] and Next3D [61] are single-shot. Figure 10 presents qualitative self-reenactment results, with additional quantitative LPIPS scores: Ours (0.0236), InvertAvatar (0.0962), Portrait4D (0.0843), and Next3D (0.2274). Both results show that SynShot significantly outperforms the baselines. Moreover, Figure 8 presents expression transfer, where our method best captures the subject’s identity and is significantly more stable for novel views and expressions, whereas GAN-based methods tend to introduce substantial artifacts, particularly in side views.
@@
Table 1 presents an ablation study of our VQ-VAE architecture. Each model was evaluated on 50 test actors excluded from the training set. Our best model, in terms of sharpness and quality, regresses a feature map $\hat{\boldsymbol{x}}_{\mathrm{feat}}\in\mathbb{R}^{H\times W\times F}$, where $F=128$, at a resolution of $512\times 512$. Regressing Gaussian primitives directly (No Sampling) suffers from lack of quality. Using a Single Layer instead of two (for hair + face) results in a lower number of Gaussians, which also decreases the final quality. A key feature of our network is densification through texture sampling. In the (Tex. up-sampling) experiment, we predict feature maps at $256\times 256$ resolution compared to $512\times 512$ and apply bilinear sampling to upscale the per-region sampled feature maps to $512\times 512$. This approach achieves results that are almost on par while saving VQ-VAE computation and memory as it operates in a downsampled texture space. Finally, using codebook quantization of latent space improves the final image quality (w/o VQ).

### v3

@@ 1 Introduction
Another recent body of work addresses the problem of building 3D head avatars from one or few input images, [77, 8, 11]. However, their rendering quality and fidelity are typically lower than those of methods trained on large datasets (e.g., [50, 88]). To improve quality, some methods [81, 75, 76] first learn a multi-identity head model that is used as prior when optimizing for the personalized avatar. Training these head priors requires a large-scale multi-view image dataset that is expensive and time-consuming to capture. Moreover, managing real data under protection laws like GDPR is cumbersome for experimentation and maintenance, as users must periodically (e.g., every 30 days) delete all dataset derivatives and trained models, allowing dataset participants to be removed from both if needed. Alternatively, the FFHQ dataset [29] may be employed, with 4D GAN-based methods [80, 61, 13] constructing an inversion prior from it. However, these approaches tend to exhibit artifacts during novel view synthesis and struggle with preserving identity. In summary, the expressive power of this prior is strongly influenced by: the training data diversity (e.g., ethnicity, age, facial features, expressions), the multi-view capture hardware setup (i.e., lighting, view-density, calibration quality, frame-rate), and the quality of the data pre-processing (e.g., mesh tracking, background masking).
@@
In contrast to the previous work that focuses on expensive and cumbersome real data, we overcome these limitations and propose SynShot, a new method that builds a prior solely on synthetic data and adapts to a real test subject requiring only a few input images. Building on the success of ML models trained on synthetic data for tasks like 3D face regression [57], 2D landmark prediction [70], rigid face alignment [3], and few-shot head reconstruction [68, 78, 6], SynShot is trained solely on a large synthetic dataset generated from 3DMM samples and diverse assets. Synthetic data offers complete control over dataset creation to meet size and diversity needs for training an expressive head prior, eliminating the need for costly capture hardware and addressing privacy concerns with real subjects. The benefits brought by synthetic data come at the cost of having to handle the domain gap between the trained head prior and real images captured “in the wild”. To effectively bridge this gap, we first fit the synthetic prior to real images and then fine-tune the prior weights to the real data using the pivotal tuning strategy proposed in [52]. With as few as three input images, SynShot reconstructs a photorealistic head avatar that generalizes to novel expressions and camera viewpoints (Fig. 1). The results show that our method outperforms state-of-the-art personalized monocular methods [85, 71, 59] trained on thousands of images each. Our method represents head avatars using 3D Gaussian primitives [30], where Gaussian parameters are generated by a convolutional architecture in UV space [41, 88, 55].
@@
    A generative method based on a convolutional encoder-decoder architecture that is trained on extensive synthetic data only to produce controllable 3D head avatars.
@@ 2 Related Work
Recent methods by Xu et al. [76] are conceptually similar to HeadNeRF and MofaNeRF; however, instead of embedding NeRF [45] on a mesh, they employ 3DGS [30]. GPHM [76] uses a series of MLPs to generate Gaussian primitives attached to a parametric model, enabling expression control and inversion, though it conditions only the avatar’s shape. GPHMv2 [75] extends GPHM with a dynamic module for improved reenactment control and a larger dataset, further enhancing quality. HeadGAP [81] also models avatars using MLPs, utilizing part-based features and additional color conditioning to improve quality. While these methods embed primitives directly on the mesh surface, our approach explicitly learns the primitive parameters by modeling their distribution via a VQ-VAE [66], eliminating the need for a guiding mesh during the test time as the shape is captured within our latent space.
@@
Volumetric primitives, combined with multi-view training, are highly effective for modeling human heads [88, 55, 74, 23, 32, 28, 63, 62] as they capture intricate details like hair and subsurface scattering [56].VolTeMorph [21] embeds a NeRF within tetrahedral cages that guide volumetric deformation. Qian et al. [50] attach Gaussian primitives to 3DMM triangles, whose local rotations and stretch deform the Gaussians without requiring neural networks. Xu et al. [74] and Giebenhain et al. [23] extend that work to further predict corrective fields over the Gaussians; rather than colors, they splat features that are translated into color by an image-space CNN [65]. Lombardi et al. [42] position 3D voxels with RGB and opacity values at the vertices of a head mesh, using ray tracing for volumetric integration. Saito et al. [55] improve quality by replacing voxel primitives with 3D Gaussians and applying rasterization. Our VQ-GAN training aligns with these principles for few-shot capture, as we supervise the process using a hybrid mesh-primitive approach to model the generative distribution.
@@
Monocular methods often rely on a strong 3DMM prior, as recovering a 3D shape from a 2D image is an inherently under-constrained problem. Face2Face [64] was a seminal work that enabled real-time reconstruction and animation of a parametric model. However, it lacks detailed hair representation and relies on low-frequency PCA texture models, which significantly affects quality. This limitation has led to the rise of neural avatars based on NeRF [67, 77, 73, 82, 72, 10, 82, 83, 18, 20, 19] and later on 3D Gaussian primitives [71, 59, 33, 9]. INSTA [85] applies triangle deformation gradients [60] to each NeRF sample based on proximity to the nearest triangle, enabling avatar animation. This approach has been adapted to 3D Gaussian Splatting (3DGS) by methods like Flash Avatar [71] or Splatting Avatar [59]. Unlike SynShot, these monocular methods do not generalize well to novel views and expressions. Moreover, they require three orders of magnitude more real data to create a single avatar. SynShot overcomes this by leveraging a synthetic prior during few shot avatar inversion, achieving high-quality results.
@@ 3 Method
This section describes SynShot, how we train the synthetic prior to generate drivable 3D Gaussian head avatars, and how we use it for few-shot head avatar reconstruction.
@@ 3.1 Preliminaries
We represent a base 3D mesh as $\mathbf{S}=\bar{\mathbf{S}}+\boldsymbol{\delta}\mathbf{B}_{id}+\boldsymbol{\gamma}\mathbf{B}_{expr}$, where $\bar{\mathbf{S}}$ is the average shape, $\mathbf{B}_{(id,expr)}$ are the bases for identity and expression of a 3DMM, and $\boldsymbol{\delta}$, $\boldsymbol{\gamma}$ denote the corresponding coefficients. Additionally, we use linear blend skinning (LBS) for head rotation around the neck with pose corrective offsets, and to rotate the eyeballs.
The head avatar is rendered via 3D Gaussian splatting, using the CUDA implementation of 3DGS [30]. The rasterizer is defined as $\mathcal{R}(\boldsymbol{G},\mathbf{K})\rightarrow\bar{\mathcal{I}}$, for a camera $\mathbf{K}$ and a set of $n$ 3D Gaussians $\boldsymbol{G}\in\mathbb{R}^{n\times(11+16\times 3)}:=\{\boldsymbol{\phi},\boldsymbol{\theta},\boldsymbol{\sigma},\boldsymbol{\alpha},\boldsymbol{h}\}$, with position $\boldsymbol{\phi}\in\mathbb{R}^{n\times 3}$, rotation $\boldsymbol{\theta}\in\mathbb{R}^{n\times 3\times 3}$, scale $\boldsymbol{\sigma}\in\mathbb{R}^{n\times 3}$, opacity $\boldsymbol{\alpha}\in\mathbb{R}^{n}$, and the (third-degree) spherical harmonics parameters $\boldsymbol{h}\in\mathbb{R}^{n\times 16\times 3}$, where $n$ is the number of Gaussians. See Kerbl et al. [30] for more details.
@@ 3.2 Gaussian Prior Model
A common limitation of using CNNs to regress Gaussian maps is the fixed output resolution, which ties the number of primitives to the output dimensions. This restriction can significantly limit the quality of the reconstructed avatar (see Table 1). To address this issue, we use a part-based densification mechanism. Similar to Kirschstein et al. [33], we use bilinear sampling, $\mathcal{B}(\cdot,u,v)$ to sample the output of the decoders at UV-positions $(u,v)$. As different head regions $r\in\{face,hair\}$ have varying requirements for the density of Gaussian primitives, we bilinearly sample separate parameter maps for the face and scalp region, rather than a single joint map. Thus, per-part map sampling acts as adaptive primitive densification for the individual regions to improve visual quality (Table 1).
@@ 3.3 Few-shot Avatar Reconstruction
To make the problem tractable, we employ a few heuristics to aid the optimization. These include early stopping with a warmup phase and an exponential moving average on the loss to determine the stopping criteria. Additionally, we scale the number of optimization steps based on the number of training frames, using a constant factor of 10 to increase the likelihood that each sample is seen at least once. As a training objective, in addition to our photometric term $\mathcal{L}_{\text{color}}$ (Eq. 5), we follow Lattas et al. [39] and, based on ArcFace [12], define two additional objectives: $\mathcal{L}_{\text{id}}$ and $\mathcal{L}_{\text{arc}}$. The final inversion loss is equal to $\mathcal{L}=\mathcal{L}_{color}+\mathcal{L}_{arc}+\mathcal{L}_{id}$. For a number of views, up to 20, the optimization takes less than 10 minutes on a single Nvidia H100 which is comparable to INSTA [85]. The training time increases with the number of frames as we scale the iterations accordingly.
@@ 3.4 Synthetic Dataset
Our dataset consists of approximately 2,000 unique identities, which we render with resolution $768\times 512$ using Blender (Cycles) following Wood et al. [70], see Fig. 4. We positioned fourteen cameras in front of the subject and an additional fourteen cameras sampled from the upper hemisphere, centered on the scene. We randomly assign assets such as hairstyles and beards to these avatars. Additionally, we utilize high-quality face textures which are randomly distributed among the samples. By combining different shapes and appearances, we augment the set of identities, following practices in synthetic data [70] and 3D face reconstruction [39, 46, 14]. To incorporate tracked expressions from multi-view setups, we propagate them to the avatars during sequence rendering. We additionally compute a hair proxy from strands by voxelizing and fitting it to the scalp region; we apply the same approach for beards. Using a neutral mesh and its hair proxy, we backproject the images onto the texture map. During test time, we use a 3DMM regressor and the input images to extract a texture, which is then used as an initialization for our method. In total, our dataset comprises 14 million images.
@@
Figure 6: Novel view evaluation of long hair and beard inversion using only three input images demonstrates the strong generalization capability of SynShot.
Figure 8: The GAN-based self-reenactment comparison again shows that SynShot better captures identity and synthesizes novel views, proving its usefulness as a synthetic prior and for pivotal fine-tuning in inversion. LPIPS scores: Ours (0.0236), InvertAvatar (0.0962), Portrait4D (0.0843), and Next3D (0.2274). Inversion input images can be found in the supplemental materials.
Figure 9: The GAN-based cross-reenactment comparison shows that our method better reconstructs the target subject’s appearance (identity) and remains faithful to the source subject’s head poses and expressions, whereas the other methods suffer from artifacts.
@@ 4 Results
We compare SynShot to two different types of methods, state-of-the-art personalized monocular methods, and inversion-based general methods. The personalized monocular methods are controlled by FLAME [40] meshes and include INSTA [85], Flash Avatar [71], and SplattingAvatar [59]. For monocular methods, we used an ensemble of four datasets [25, 88, 82, 19] processed using the face tracker from Zielonka et al. [86]. SplattingAvatar follows the approach of Zheng et al. [82] and uses the monocular 3D face regressor DECA [17] for tracking. In our experiments, we adopted a similar approach, employing an in-house regressor, similar to DECA, to estimate 3DMM expression and pose parameters. While these methods produce photorealistic avatars, they struggle with generalization to novel views and poses (see Figure 5). For inversion-based methods, we compare PanoHead [2], HeadNeRF [26], and MofaNeRF [84]. We also compare to concurrent works including Portrait4D [13], Next3D [61], and InvertAvatar [80] (Figures 8 and 9). We use three images for all inversion experiments see supp. material. Figure 6 presents a novel view evaluation of challenging long hair and beard inversion, demonstrating the generalization capabilities of SynShot.
@@
To measure the performance of SynShot without introducing bias, we selected training frames from $\{F_{n}\}_{n=1}^{16}=\{1,\dots,987\}$, where $F_{n}$ denotes the Fibonacci sequence. For all experiments, we use progressive farthest point sampling [49] in the 3DMM expression space to select a specified number of frames from the training set. The self-reenactment sequences were evaluated using LPIPS and SSIM on the last 600 frames from the INSTA dataset [85].
@@
Our combined dataset consists of eleven monocular sequences ($512\times 512$ resolution), many of which are in-the-wild videos with very limited head motion, resulting in a low error as the test sequences closely resemble the training data, leaving limited room to assess diversity. To address this and accurately measure the effective error, we trained each method on a varying number of frames, corresponding to frames used in our inversion pipeline. The reconstruction error is evaluated on 600 test frames. Figure 7 demonstrates the effectiveness of our inversion, particularly with up to 233 training frames. Due to the lack of a strong prior, monocular methods fail in low training frame regimes, and, even with larger training datasets, they do not perform well and produce artifacts. Please note that to benefit from an increased number of input frames, i.e., to ground the avatar reconstruction more on the input than the synthetic prior, it requires an increased number of optimization iterations during pivotal fine-tuning. The number of iteration steps affects the metrics, causing LPIPS to vary non-monotonically.
@@
We compared SynShot to three animatable GAN-based methods. For our method and InvertAvatar [80], we used three input images, whereas Portrait4D [13] and Next3D [61] are single-shot. Figure 8 presents qualitative self-reenactment results, with additional quantitative LPIPS scores: Ours (0.0236), InvertAvatar (0.0962), Portrait4D (0.0843), and Next3D (0.2274). Both results show that SynShot significantly outperforms the baselines. Moreover, Figure 9 presents expression transfer, where our method best captures the subject’s identity and is more stable for novel views and expressions, whereas GAN-based methods tend to introduce artifacts in side views.
@@
Table 1 presents an ablation study of our VQ-VAE architecture. Each model was evaluated on 50 test actors excluded from the training set. Our best model, in terms of sharpness and quality, regresses a feature map $\hat{\boldsymbol{x}}_{\mathrm{feat}}\in\mathbb{R}^{H\times W\times F}$, where $F=128$, at a resolution of $512\times 512$. Regressing Gaussian primitives directly (No Sampling) suffers from lack of quality. Using a Single Layer instead of two (for hair + face) results in a lower number of Gaussians, which also decreases the final quality. A key feature of our network is densification through texture sampling. In the (Tex. up-sampling) experiment, we predict feature maps at $256\times 256$ resolution compared to $512\times 512$ and apply bilinear sampling to upscale the per-region sampled feature maps to $512\times 512$. This approach achieves results that are almost on par while saving VQ-VAE computation and memory.Finally, using codebook quantization of latent space improves the final image quality (w/o VQ).